### v3

# Jin Li    Ye Luo    Xiaowei Zhang

Seesaw Experimentation: A/B Tests with Spillovers thanks: Contact information: Jin Li (jli1@hku.hk), Ye Luo (kurtluo@hku.hk), Faculty of Business and Economics, The University of Hong Kong, Pokfulam Road, Hong Kong SAR. Xiaowei Zhang (xiaoweiz@ust.hk), Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong SAR. We thank Shan Huang and Ruohan Zhan for helpful discussions.
Jin Li    Ye Luo    Xiaowei Zhang
(December 31, 2024)
This paper investigates how a firm’s performance can decline even when consistently implementing innovations validated through A/B testing. We introduce the concept of seesaw experimentation, where successful innovations enhance performance in measured primary metrics but generate negative spillover effects in unmeasured secondary dimensions, ultimately reducing overall performance. We identify the conditions under which seesaw experimentation occurs. We also propose a simple solution to address it: implementing a positive hurdle rate for A/B tests. We derive the optimal hurdle rate for both the normal distribution and the fat-tailed distributions.
Keywords: A/B test, experimentation, spillovers
JEL Codes: C12, D62, L25
1 Introduction
A/B testing has become an important tool for modern business, enabling firms to drive innovation, improve user experiences, and increase revenue growth (Kohavi et al., 2020, Koning et al., 2022). Major tech firms such as Amazon, Google, Meta, and Microsoft deploy thousands of such tests annually, embedding data-driven decision making into their innovation processes (Thomke, 2020, Azevedo et al., 2020).
While existing research typically examines A/B tests in isolation, these experiments frequently generate significant spillover effects [^1] We use spillover, rather than negative spillover, for expositional ease. that ripple across an organization’s ecosystem of products, channels, and performance metrics. These effects are particularly pronounced on large digital platforms. At the WeChat platform, for instance, enhancing Video Accounts can substantially decrease engagement with Official Accounts as users redistribute their limited attention. The spillover effects extends beyond digital interactions: improvements that drive online sales growth often cannibalize offline revenue, and strategies that maximize advertising revenue frequently degrade user experience. These examples underscore a fundamental challenge: improvements in one area of business operations can trigger unintended consequences in seemingly unrelated domains. As externalities are prevalent within organizations, so are the spillover effects.
These spillover effects create what we term seesaw experimentation, where the overall performance of a firm deteriorates despite continuous improvements through innovations in current strategic priorities. The name reflects its oscillating nature. Like a seesaw, advances in one dimension often diminish performance in others. Firms implement innovations that boost their primary metrics—whether in specific products, channels, or KPIs. Yet these same innovations generate hidden or underweighted losses in secondary dimensions. Because business environments and strategic priorities constantly evolve, yesterday’s focus on gross merchandise value may shift to user satisfaction tomorrow. Each shift triggers new waves of innovations that may inadvertently reverse previous gains. The pattern creates a dangerous illusion of progress. Success metrics in current focal areas shine brightly. Meanwhile, the firm’s overall performance quietly degrades.
Managing spillover effects in A/B testing poses a formidable challenge. This is especially true for large organizations running numerous simultaneous experiments. Some firms try to address single-product spillovers through managerial coordination. For instance, user growth teams may challenge revenue-boosting innovations that hurt customer satisfaction. But such efforts often fall short. Several factors drive this coordination failure. First, many potential effects—especially long-term impacts—defy easy quantification. Second, the sheer volume of concurrent experiments makes comprehensive coordination impractical. Third, cross-product and cross-channel effects often remain hidden from view. Despite widespread recognition of their importance for platform health, spillover effects rarely factor into experimental design [^2] Communications with experts at a leading Chinese platform. This creates a persistent gap between the ideal of holistic testing and the reality of fragmented experimentation.
The purpose of this paper is to study A/B testing in the presence of spillover effects, with particular attention to the emergence and implications of seesaw experimentation. We investigate the conditions under which firms, despite achieving continuous improvements in primary performance metrics, may experience deteriorating overall outcomes. Our analysis focuses on two key questions: First, under what conditions does seesaw experimentation emerge as firms improve their primary strategic objectives? Second, how can organizations effectively manage these spillover effects without relying on complex cross-managerial coordination?
We develop a bivariate normal distribution framework to demonstrate the emergence of seesaw experimentation. Our model identifies conditions where firms achieve consistent A/B test successes yet face declining overall performance. Two key factors drive this phenomenon. First, a higher signal-to-noise ratio (the absolute value of mean over standard deviation) increases the likelihood of seesaw experimentation. Second, this effect strengthens as the correlation between dimensions becomes more negative.
Our analysis yields a simple yet powerful solution to manage spillover: implement a positive hurdle rate for A/B test approval. This approach only implements innovations that exceed a specified performance threshold. We derive the optimal hurdle rate where the positive gains in the primary dimension equal the negative externalities in the secondary dimension. Our framework provides practical guidance by showing how this optimal rate varies with key distribution parameters. This solution effectively balances innovation benefits against spillover costs without requiring complex coordination mechanisms.
The literature addressing our specific focus is relatively small. Berman et al., (2018) examine how p-hacking in A/B testing can lead to false discoveries and impair experimental improvements. Azevedo et al., (2020) analyze how distributional characteristics, particularly fat tails, influence experimentation allocation strategies. [^3] We extend our analysis to fat-tailed distributions in Section 4. McClellan, (2022) explores how agency problems affect A/B test adoption mechanisms. While these studies provide valuable insights, they primarily focus on single-dimensional outcomes. Our work departs from this literature by explicitly modeling the interconnectedness of performance across multiple dimensions, revealing how improvements in one dimension can generate spillover effects in others.
2 Model Formulation
Consider a firm whose management is concerned with two performance dimensions, denoted as $u$ and $v$. In each period $t=1,\ldots,T$, the firm identifies its strategic priority $a_{t}\in\{u,v\}$ between these two dimensions. We refer to this prioritized dimension as the primary dimension, and the other as the secondary dimension.
The firm then conducts an A/B test to evaluate a potential innovation using a metric aligned with the primary dimension $a_{t}$, measuring the innovation’s effect in this dimension. Let $u_{t}$ and $v_{t}$ represent the innovation’s effects on dimensions $u$ and $v$, respectively. Note that only the effect on the primary dimension is measured in the A/B test. [^4] This assumption reflects practical challenges: initiators of the A/B test may not have the incentive to measure the outcomes in secondary dimensions, and, in general, outcomes are often harder to measure and may rely on private information held by different stakeholders. But even if firms can perfectly measure outcomes from the secondary dimension, their decisions depend on how to weigh the relative importance of the two. Our analysis reflects the limiting case where the importance of the primary dimension (at the moment of the decision) far outweighs the secondary dimension.
Subsequently, the firm decides whether to adopt this innovation based on the measured effect. Let $d_{t}\in\{0,1\}$ denote the adoption decision in period $t$, where $1$ indicates adoption and 00 indicates non-adoption. The decision rule can be expressed as:
$d_{t}=\mathbb{I}(a_{t}=u,u_{t}>0)+\mathbb{I}(a_{t}=v,v_{t}>0).$
In other words, the firm adopts the innovation if and only if it improves the firm’s performance in the primary dimension $a_{t}$ for the current period $t$.
In this paper, we examine the firm’s cumulative performance through a multi-period innovation process driven by A/B testing. Let
$U_{T}=\sum_{t=1}^{T}d_{t}u_{t}\quad\mbox{and}\quad V_{T}=\sum_{t=1}^{T}d_{t}v_{t}.$
Here, $U_{T}$ represents the cumulative performance in dimension $u$ of the adopted innovations over the time horizon $T$, while $V_{T}$ represents the same for dimension $v$. We define the firm’s overall performance at time $T$ as the sum of these cumulative performances: $U_{T}+V_{T}$.
Suppose that the market environment changes exogenously and management accurately identifies strategic priorities to adapt over time. We model this by treating the strategic priorities $a_{t}$ as independent and identically distributed (i.i.d.). Furthermore, suppose that each potential innovation’s effects $(u_{t},v_{t})$ on both dimensions are i.i.d. across time periods, while allowing for correlation between dimensions within each period.
Formally, we make the following assumption throughout the paper:
Assumption 1.
  1. (i)
    $\{a_{t}:t\geq 1\}$ is a sequence of i.i.d. realizations of a random variable $A$, where $\mathbb{P}(A=u)=p_{u}$, $\mathbb{P}(A=v)=p_{v}$, and $p_{u}+p_{v}=1$. $\{(u_{t},v_{t}):t\geq 1\}$ is a sequence of i.i.d. realizations of a random vector $(U,V)$.
  2. (ii)
    The sequences $\{a_{t}:t\geq 1\}$ and $\{(u_{t},v_{t}):t\geq 1\}$ are independent of each other.
Under Assumption 1, the firm’s overall performance is a sum of i.i.d. random variables:
$U_{T}+V_{T}=\sum_{t=1}^{T}[d_{t}(u_{t}+v_{t})].$
The strong law of large numbers then implies that
$\mathbb{P}\left(\lim_{T\to\infty}\frac{U_{T}+V_{T}}{T}=\mathbb{E}[D(U+V)]\right)=1.$
Definition 1.
A firm exhibits the phenomenon of seesaw experimentation when its long-run average performance $\mathbb{E}[D(U+V)]$ is negative despite adopting innovations only when they demonstrate positive effects in their respective primary dimensions.
For simplicity, we start by considering bivariate normal distributions. We extend our results to a wider class of distributions later.
Assumption 2.
  1. (i)
    $(U,V)$ follows a bivariate normal distribution with mean vector $m$ and covariance matrix $\Sigma$:
    $m=\begin{pmatrix}\mu_{u}\\ \mu_{v}\end{pmatrix}\quad\mbox{and}\quad\Sigma=\begin{pmatrix}\sigma_{u}^{2}&\rho\sigma_{u}\sigma_{v}\\ \rho\sigma_{u}\sigma_{v}&\sigma_{v}^{2}\end{pmatrix}.$
    Here, $\sigma_{u}^{2}$ and $\sigma_{v}^{2}$ are the marginal variances of $U$ and $V$, respectively, and $\rho$ is their correlation.
  2. (ii)
    Both $\mu_{u}$ and $\mu_{v}$ are negative.
Assumptions 1 and 2 allow us to deliver the main message of the paper—the existence of seesaw experimentation and its mitigation via a positive hurdle rate—in the simplest and most standard framework. The conditions in these assumptions can be relaxed without affecting the main points of the paper. In particular, in Assumption 1(i), $a_{t}$ can be correlated over time, and Assumption 1(ii) is also not necessary in the sense that our main results hold even if the choice of innovation can depend on arbitrary realizations of outcomes in the past. For Assumption 2(i), our results also apply to broader classes of distributions, including those having fat tails and those whose values can on discrete values. Assumption 2(ii) is not necessary in the sense that seesaw experimentation can still occur in scenarios where either $\mu_{u}$ or $\mu_{v}$ is non-negative while the other is negative, provided that $\mu_{u}+\mu_{v}<0$. But we assume this condition to reflect the fundamental business reality that random innovations typically do not improve secondary performance dimensions. Below, we characterize conditions under which seesaw experimentation occurs and analyze how to determine an optimal hurdle rate to address this issue.
3 Base Model
We first consider the scenario where the two dimensions are symmetric. In other words, while they can be correlated, $U$ and $V$ have the same marginal distribution: $\mu_{u}=\mu_{v}$ and $\sigma_{u}=\sigma_{v}$. We use $\mu$ and $\sigma$ to represent their common values, respectively.
Using the law of iterated expectations and the independence between $A$ and $(U,V)$, we can express $\mathbb{E}[D(U+V)]$ as:
$\mathbb{E}[D(U+V)]$
[…]
Hence, a sufficient condition for $\mathbb{E}[D(U+V)]$ to be negative is that both $\mathbb{E}[U+V|U>0]$ and $\mathbb{E}[U+V|V>0]$ are negative. These quantities can be calculated explicitly under the bivariate normal assumption.
Proposition 1.
Suppose Assumptions 1 and 2 hold and $(U,V)$ follows a symmetric distribution. If
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)M\left(\frac{-\mu}{\sigma}\right)-1,$ (1)
then $\mathbb{E}[D(U+V)]<0$, where $M(\alpha)\coloneqq(1-\Phi(\alpha))/\phi(\alpha)$ is the Mills ratio of the standard normal distribution, with $\phi$ and $\Phi$ representing the probability density function and cumulative distribution function of the standard normal distribution, respectively.
Proposition 1 shows that seesaw experimentation can occur. That is, while a firm may seemingly improve its performance constantly through successful A/B tests, its overall performance may deteriorate over time. This decline stems from significant negative spillover effects on secondary dimensions that offset the gain in the primary dimension.
Proposition 1 also provides a sufficient condition for seesaw experimentation to occur. One factor is the signal-to-noise ratio $\alpha=|\mu|/\sigma$. We show in the Appendix (Lemma 1) that $\alpha M(\alpha)$ is an increasing function of $\alpha$, implying that the sufficient condition is more likely to be satisfied when signal-to-noise ratio increases. This can happen in two ways: either through an increase in $|\mu|$ or a decrease in $\sigma$. The intuition behind this is straightforward. When $|\mu|$ is larger, the negative impact on the secondary dimension becomes more pronounced. Alternatively, when $\sigma$ is smaller, there’s less variability in outcomes, which typically results in smaller average benefits in the primary dimension. Both scenarios make seesaw experimentation more likely to occur.
Another key factor is the correlation coefficient $\rho$ between the two dimensions. As $\rho$ becomes more negative, seesaw experimentation becomes more likely, reflecting situations where improvements in the primary dimension are increasingly linked to negative effects in the secondary dimension. While one might intuitively expect seesaw experimentation to occur only with negative correlation, our analysis reveals that this is not true. Because $(2\alpha M(\alpha)-1)$ monotonically increases from $-1$ to $1$ for $\alpha\in[0,\infty)$, our sufficient condition indicates that seesaw experimentation can emerge even with positive correlation. This counterintuitive result occurs, for example, when $|\mu|$ is large, meaning the average negative effect on the secondary dimension is sufficiently strong to override the positive correlation between dimensions.
How to deal with seesaw experimentation? The traditional approach relies on inter-departmental coordination. For instance, customer satisfaction executives might veto an A/B test that boosts revenue but diminishes customer experience. However, this approach faces several challenges. The relative importance of different dimensions often involves subjective judgment, and outcomes can vary depending on individual executives’ perspectives and personalities. Moreover, when companies carry out thousands of A/B tests, the sheer volume of experiments and the difficulty in measuring secondary dimension effects make coordination costs prohibitively high.
Below, we propose a coordination-free solution: the implementation of a positive hurdle rate $z$. This approach requires that any innovation must demonstrate a benefit exceeding $z$ in the primary dimension before adoption, effectively raising the bar for implementation without requiring complex coordination across departments.
When a positive hurdle $z$ is implemented, the firm’s long-run average performance becomes:
$\mathbb{E}[D(U+V)]={}$ $\mathbb{P}(A=u)\mathbb{P}(U>z)\mathbb{E}[U+V|U>z]+\mathbb{P}(A=v)\mathbb{P}(V>z)\mathbb{E}[U+V|V>z]$
[…]
Proposition 2.
Suppose Assumptions 1 and 2 hold and $(U,V)$ follows a symmetric distribution. If $\rho\in(-1,1)$, then $f^{\prime}(0)>0$, $f(z)$ attains its maximum at $z^{*}$ and $f(z^{*})>0$, where
$z^{*}=\frac{\rho-1}{\rho+1}\mu.$
Proposition 2 demonstrates that implementing a positive hurdle rate is strictly beneficial to the firm. The intuition is straightforward: When the firm sets a hurdle rate slightly above zero, there are two conflicting effects. First, there is a small loss in the primary dimension because some marginally beneficial innovations are not implemented. Note that the foregone benefits are small for these innovations. Second, there is a significant gain in the secondary dimension because the firm avoids implementing innovations that would have negative effects (in expectation) there. By setting a positive hurdle rate, the firm screens out potential innovations that might have small positive benefits in the primary dimension but large negative effects in the secondary dimension. This makes a positive hurdle rate beneficial.
This proposition also determines the optimal hurdle rate and reveals two key relationships. First, the optimal hurdle rate is proportionally related to $\mu$ (the mean value in the secondary dimension). When $\mu$ is lower, the firm sets a higher hurdle rate. This makes intuitive sense, as the firm wants to be more selective when secondary outcomes are likely to be less favorable.
Second, the proposition links the optimal hurdle rate with the correlation $\rho$ between primary and secondary outcomes. Note that when $\rho=0$ (no correlation), the optimal hurdle rate equals $-\mu$, which is the expected negative impact in the secondary dimension. This arises because, with no correlation, $-\mu$ is the average negative secondary effect. When $\rho$ approaches $1$ (perfect positive correlation), the optimal hurdle rate approaches zero. This is because with perfect positive correlation, any innovation that is good for the primary dimension will also be good for the secondary dimension, eliminating the need for a positive hurdle. Finally, when $\rho$ approaches $-1$ (perfect negative correlation), the optimal hurdle rate approaches infinity. This extreme case occurs because perfect negative correlation means that any positive outcome in the primary dimension necessarily creates an equally large negative outcome in the secondary dimension, making the overall gain from innovation small.
The optimal hurdle rate is set by a basic economic principle: it should be where the positive gain from an innovation in the primary dimension equals the negative externality on the secondary dimension. At the optimal hurdle rate, the externality is internalized. Under the symmetric bivariate normal assumption, the negative externality is $\mathbb{E}[-V|U=u]=-\rho u-(1-\rho)\mu$. As shown in Figure 1, the gain in the primary dimension exceeds the expected negative externality only if it surpasses the optimal threshold $z^{*}=\frac{1-\rho}{1+\rho}\mu$.
$u$$v$$v=u$$v=\mathbb{E}[-V|U=u]$$z^{*}=\frac{\rho-1}{\rho+1}\mu$
Figure 1: Negative Externality and Optimal Hurdle Rate
Finally, note that the optimal hurdle rate in Proposition 2 maximizes the firm’s long-term performance. A firm that also prioritizes short-term objectives should set a lower hurdle rate, accepting more innovations despite their potential negative secondary effects. This reflects the classic trade-off between short-term gains and long-term losses.
4 Extensions
This section extends the results of Section 3 in three directions: asymmetry, multidimensionality, and fat tails. For clarity of managerial insights, we examine each extension separately rather than analyzing a fully general model combining all features.
4.1 Asymmetry
Consider a general bivariate normal distribution without the symmetry assumptions $\mu_{u}=\mu_{v}$ and $\sigma_{u}=\sigma_{v}$. Due to this asymmetry between dimensions, the firm sets distinct hurdle rates $z_{u}$ and $z_{v}$. The firm’s long-run average performance then becomes:
$\mathbb{E}[D(U+V)]$
$={}\mathbb{P}(A=u)\mathbb{P}(U>z_{u})\mathbb{E}[U+V|U>z_{u}]+\mathbb{P}(A=v)\mathbb{P}(V>z_{v})\mathbb{E}[U+V|V>z_{v}]$
$\coloneqq g(z_{u},z_{v}).$ (3)
Proposition 3.
Suppose Assumptions 1 and 2 hold. If $-1\leq\rho<\min(\rho_{1},\rho_{2})$, then $g(0,0)<0$, where
$\rho_{1}={}$ $\left[\left(\frac{-\mu_{u}}{\sigma_{u}}\right)M\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\left(1+\frac{\mu_{v}}{\mu_{u}}\right)-1\right]\frac{\sigma_{u}}{\sigma_{v}},$
$\rho_{2}={}$ $\left[\left(\frac{-\mu_{v}}{\sigma_{v}}\right)M\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\left(1+\frac{\mu_{u}}{\mu_{v}}\right)-1\right]\frac{\sigma_{v}}{\sigma_{u}}.$
Proposition 4.
Suppose Assumptions 1 and 2 hold. If
$-\min\left(\frac{\sigma_{v}}{\sigma_{u}},\frac{\sigma_{u}}{\sigma_{v}}\right)<\rho<\min\left(\frac{\mu_{v}/\sigma_{v}}{\mu_{u}/\sigma_{u}},\frac{\mu_{u}/\sigma_{u}}{\mu_{v}/\sigma_{v}}\right),$
then $\nabla g(0,0)>0$, $g(z_{u},z_{v})$ attains its maximum at $(z_{u}^{*},z_{v}^{*})$, and $g(z_{u}^{*},z_{v}^{*})>0$ where
$z_{u}^{*}=\frac{\rho\mu_{u}/\sigma_{u}-\mu_{v}/\sigma_{v}}{\rho/\sigma_{u}+1/\sigma_{v}}\quad\mbox{and}\quad z_{v}^{*}=\frac{\rho\mu_{v}/\sigma_{v}-\mu_{u}/\sigma_{u}}{\rho/\sigma_{v}+1/\sigma_{u}}.$
The conditions in Proposition 3 reduce to that in Proposition 1 when the means and variances of the two dimensions are equal. Proposition 4 derives the optimal hurdle rates for both dimensions. The optimal hurdle rate for each dimension follows the same principle as in the symmetric case: it is set to the value at which the positive gain from adopting an innovation exactly offsets its negative externality on the other dimension.
The optimal hurdle rate increases when either the mean of the secondary dimension is lower or when the correlation coefficient is lower, following the same logic as in the symmetric case since both conditions lead to greater negative externalities.
In addition, the optimal hurdle rate also depends on the mean of the primary dimension, except when the correlation coefficient $\rho$ is zero. In the special case where $\rho=0$, the hurdle rate equals $-\mu$ (the negative externality imposed on the other dimension). When $\rho$ is negative, the optimal hurdle rate $z_{u}^{*}$ decreases as $\mu_{u}$ increases, because a higher mean in the primary dimension reduces the negative externality, leading to a higher optimal hurdle rate. [^5] Under the bivariate normal assumption, the negative externality is $\mathbb{E}[-V|U]=\rho(\sigma_{v}/\sigma_{u})(\mu_{u}-U)-\mu_{v}$. As a result, when $\rho<0$, an increase in $\mu_{u}$ reduces the negative externality. Similarly, when $\rho$ is positive, $z_{u}^{*}$ increases in $\mu_{u}$.
4.2 Multidimensionality
Now suppose the firm measures performance across $n$ dimensions. Each period, it identifies a dimension $A$ as its strategic priority, evaluates a potential innovation designed for dimension $A$ through an A/B test, and adopts the innovation if its effect on dimension $A$ exceeds a threshold $z$. Let $X_{i}$ denote the innovation’s effect on dimension $i$ ($i=1,\ldots,n$), although only the effect on the primary dimension $A$ is measured. Define the adoption decision as $D=\mathbb{I}(X_{A}>z)$. The long-run average of the firm’s overall performance is then:
$\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\right]=\sum_{i=1}^{n}\mathbb{P}(A=i)\mathbb{P}(X_{j}>z)\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\bigg{|}X_{j}>z\right]\coloneqq h(z).$ (4)
Assumption 3.
  1. (i)
    $(X_{1},\ldots,X_{n})$ follows a multivariate normal distribution with mean vector $m$ and covariance matrix $\Sigma$:
    $m=\begin{pmatrix}\mu\\ \mu\\ \vdots\\ \mu\end{pmatrix}\quad\mbox{and}\quad\Sigma=\begin{pmatrix}\sigma^{2}&\rho\sigma^{2}&\cdots&\rho\sigma^{2}\\ \rho\sigma^{2}&\sigma^{2}&\cdots&\rho\sigma^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \rho\sigma^{2}&\rho\sigma^{2}&\cdots&\sigma^{2}\end{pmatrix}.$
  2. (ii)
    $\mu<0$.
Proposition 5.
Suppose Assumptions 1 and 3 hold. Then, $h(0)<0$ if
$-\frac{1}{n-1}\leq\rho<\left(\frac{n}{n-1}\right)\left(\frac{-\mu}{\sigma}\right)M\left(\frac{-\mu}{\sigma}\right)-\frac{1}{n-1}.$
Proposition 6.
Suppose Assumptions 1 and 3 hold. If $-1/(n-1)<\rho<1$, then $h^{\prime}(0)>0$, $h(z)$ attains its maximum at $z^{*}$, and $h(z^{*})>0$, where
$z^{*}=\frac{(n-1)(\rho-1)}{(n-1)\rho+1}\mu.$
When $n=2$, the expressions in Propositions 5 and 6 reduce to those in our bivariate-normal analysis. In Proposition 5, the lower bound $-1/(n-1)$ is the lowest feasible correlation coefficient in a multivariate normal distribution. This indicates that seesaw experimentation emerges when A/B test outcomes across different dimensions are sufficiently negatively correlated, consistent with our earlier findings. Proposition 6 shows that the optimal hurdle rate increases in $n$. This reflects an intuitive relationship: more dimensions create more negative externalities, leading to a higher optimal threshold for innovation adoption.
4.3 Fat Tails
Azevedo et al., (2020) found that A/B test results frequently display fat tails that cannot be adequately modeled by normal distributions. Building on this insight, we extend our analysis to bivariate $t$ distributions. This class of distributions extends bivariate normal distributions by adding a parameter that controls tail thickness. Unlike normal distributions, their tail probabilities decay polynomially, with the added parameter determining the decay rate. [^6] While Pareto distributions also model heavy tails effectively, their bivariate extensions have limited flexibility: they are not defined over the entire real plane, and their correlation structure is constrained by the tail thickness parameter rather than being independently specified (Arnold, 2015).
Assumption 4.
  1. (i)
    $(U,V)$ follows a bivariate $t$ distribution with mean vector $m$, scale matrix $\Sigma$, and degrees of freedom $\delta$:
    $m=\begin{pmatrix}\mu\\ \mu\end{pmatrix}\quad\mbox{and}\quad\Sigma=\begin{pmatrix}\sigma^{2}&\rho\sigma^{2}\\ \rho\sigma^{2}&\sigma^{2}\end{pmatrix}.$
  2. (ii)
    $\mu<0$ and $\delta>2$.
The parameter $\delta$ controls the tail thickness of the distribution, with smaller values of $\delta$ corresponding to heavier tails. When $\delta=2$, the distribution lacks finite second moments, making covariance and correlation undefined. For $\delta>2$, the scale matrix $\Sigma$ equals the covariance matrix of $(U,V)$, where $\sigma^{2}$ represents the marginal variance and $\rho$ the correlation. As $\delta\to\infty$, the distribution converges to a bivariate normal with the same mean vector and covariance matrix, eliminating the fat tails.
Proposition 7.
Suppose Assumptions 1 and 4 hold. If
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)W\left(\frac{-\mu}{\sigma}\right)-1,$ (5)
then $f(0)<0$, where
$W(\alpha)=\left(\frac{\delta-2}{\delta}\right)\left(\frac{1-T_{\delta}(\alpha;0,1)}{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}\right),$
$t_{\delta}(\alpha;\mu,\sigma^{2})$ and $T_{\delta}(\alpha;\mu,\sigma^{2})$ denote the probability density function and cumulative distribution function, respectively, of a univariate $t$ distribution with mean $\mu$, variance $\sigma^{2}$, and degrees of freedom $\delta$.
In equation (5), the maximum correlation permitting seesaw experimentation is $12\alpha W(\alpha)-$ 1, where $\alpha=-\mu/\sigma$. Figure 2 illustrates this threshold as a function of $\alpha$ for different values of $\delta$. As $\delta$ increases, this threshold increases for any given $\alpha$. This indicates that, holding marginal mean ($\mu$) and variance ($\sigma^{2}$) constant, lighter tails expand the range of correlations that permit seesaw experimentation, making such phenomenon more likely to occur. This result has an intuitive explanation: with lighter tails, experimentation offers more limited upside potential. Consequently, the negative spillover effects in the secondary dimension are more likely to outweigh the positive gains in the primary dimension.
Figure 2: Maximum Correlation Leading to Seesaw Experimentation
Remark 1.
The standard $t$ distribution with $\delta$ degrees of freedom converges to the standard normal distribution as $\delta\to\infty$. Thus,
$W(\alpha)=\left(\frac{\delta-2}{\delta}\right)\left(\frac{1-T_{\delta}(\alpha;0,1)}{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}\right)\to\frac{1-T_{\infty}(\alpha;0,1)}{t_{\infty}(\alpha;0,1)}=\frac{1-\Phi(\alpha)}{\phi(\alpha)}=M(\alpha).$
Hence, the upper bound in equation (5) for the bivariate $t$ distribution converges to the upper bound in (1) for the base model (bivariate normal distribution).
Proposition 8.
Suppose Assumptions 1 and 4 hold. If $\rho\in(-1,1)$, then $f^{\prime}(0)>0$, $f(z)$ attains its maximum at $z^{*}$ and $f(z^{*})>0$, where
$z^{*}=\frac{\rho-1}{\rho+1}\mu.$
Note that the choice of the optimal hurdle rate is the same as that in the normal distribution. The underlying intuition, illustrated in Figure 1, holds true: the optimal hurdle rate equates the expected negative externality in the secondary dimension with the gain in the primary dimension. Both bivariate normal and bivariate $t$ distributions share the same expression for the expected negative externality: $\mathbb{E}[-V|U=u]=-\rho\mu-(1-\rho)\mu$.
5 Conclusion
Our paper examines A/B testing in the presence of spillover effects. Building on a multivariate normal framework, we identify conditions that give rise to seesaw experimentation—where continuous adoption of successful innovations leads to performance decline. We show that this phenomenon is more likely to occur when the signal-to-noise ratio is large and correlations across dimensions are more negative. We also show how implementing appropriate hurdle rates can enhance firm performance, and derive optimal thresholds within our analytical framework. At the optimal threshold, the positive gain is equal to the negative externality it imposes.
While our results are theoretical in nature, they have practical implications: when innovations have multi-dimensional impacts, firms should move away from zero-threshold A/B testing. The interconnected nature of modern organizations demands a more systematic approach to experimentation—one that explicitly accounts for these cross-dimensional externalities. Setting optimal hurdle rates in practice requires careful measurement of innovation outcomes across dimensions, including their correlations. For this purpose, firms can leverage their historical A/B test data, which provides a rich source of information about both direct effects and spillovers.
Appendices
Appendix A Proofs for Bivariate Normal Distribution
Lemma 1.
$k(\alpha)\coloneqq\alpha M(\alpha)$ is strictly increasing in $\alpha\in\mathbb{R}$, $k(0)=0$, and $\lim\limits_{\alpha\to\infty}k(\alpha)=1$.
Proof of Lemma 1.
Note: $k(\alpha)=\alpha(1-\Phi(\alpha))/\phi(\alpha)$ and $\phi^{\prime}(\alpha)=(-\alpha)\phi(\alpha)$. Hence,
$k^{\prime}(\alpha)=\frac{(1-\Phi(\alpha))(1+\alpha^{2})-\alpha\phi(\alpha)}{\phi(\alpha)}.$
Because $\phi(\alpha)>0$ and $\Phi(\alpha)\in(0,1)$ for all $\alpha$, we have $k^{\prime}(\alpha)>0$ for all $\alpha\leq 0$. Moreover, it is known that the Mills ratio satisfies
$\frac{1-\Phi(\alpha)}{\alpha}>\frac{\alpha}{1+\alpha^{2}},\quad\mbox{for all }\alpha>0;$
see, e.g., Gasull and Utzet, (2014). Hence, $k^{\prime}(\alpha)>0$ for all $\alpha>0$.
Finally, the limit of $k(\alpha)$ can be calculated via L’Hôpital’s rule: ∎
Lemma 2.
Let $g(z_{u},z_{v})$ be as defined in (4.1). Then,
$g(z_{u},z_{v})={}$ $p_{u}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right]$ (A.1)
$+p_{v}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{v}-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{z_{v}-\mu_{v}}{\sigma_{v}}\right)\right].$
Proof of Lemma 2.
We first note that
$\mathbb{E}[D(U+V)]={}$ $p_{u}\mathbb{E}[U+V|U>z_{u}]\mathbb{P}(U>z_{u})+p_{v}\mathbb{E}[U+V|V>z_{v}]\mathbb{P}(V>z_{v}).$
The expression (A.1) can be derived by using the following properties of the truncated normal distribution (Johnson and Kotz, 1972, p. 112–113): for any $z$,
$\mathbb{E}[U|U>z]=\mu_{u}+\sigma_{u}\lambda\left(\frac{z-\mu_{u}}{\sigma_{u}}\right)\quad\mbox{and}\quad\mathbb{E}[V|U>z]=\mu_{v}+\rho\sigma_{v}\lambda\left(\frac{z-\mu_{u}}{\sigma_{u}}\right),$
where $\lambda(\alpha)\coloneqq 1/M(\alpha)$ is the inverse Mills ratio of the standard normal distribution. ∎
Proof of Proposition 1.
The proof follows from $f(0)=g(0,0)$ and Lemma 2 through direct calculations. ∎
Proof of Proposition 2.
Note that
$f(z)={}$ $g(z,z)=2\mu\left(1-\Phi\left(\frac{z-\mu}{\sigma}\right)\right)+\sigma(1+\rho)\phi\left(\frac{z-\mu}{\sigma}\right),$
$f^{\prime}(z)={}$ $-\frac{1}{\sigma}\phi\left(\frac{z-\mu}{\sigma}\right)\left[2\mu+(1+\rho)(z-\mu)\right],$
$f^{\prime\prime}(z)={}$ $-\frac{1}{\sigma}\phi\left(\frac{z-\mu}{\sigma}\right)\left[\left(-\frac{1}{\sigma}\right)\left(\frac{z-\mu}{\sigma}\right)\left(2\mu+(1+\rho)(z-\mu)\right)+1+\rho\right].$
Hence, if $-1<\rho<1$, then $f^{\prime}(0)>0$ and $f^{\prime}(z)=0$ has a unique solution
$z^{*}=\frac{\rho-1}{\rho+1}\mu;$
moreover,
$f^{\prime\prime}(z^{*})=-\frac{1}{\sigma}\phi\left(\frac{-2\mu}{\sigma(1+\rho)}\right)(1+\rho)<0,$
which implies $z^{*}$ the a maximizer of $f(z)$. Moreover,
$f(z^{*})={}$ $\sigma(1+\rho)\phi\left(\frac{z^{*}-\mu}{\sigma}\right)\left[\frac{2\mu}{\sigma(1+\rho)}M\left(\frac{z^{*}-\mu}{\sigma}\right)+1\right]$
[…]
where the positivity follows from $\alpha M(\alpha)<1$ by Lemma 1. ∎
Proof of Proposition 3.
By equation (A.1), a sufficient condition for $g(0,0)<0$ is
$\left\{\begin{array}[]{l}\displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0,\\[10.00002pt] \displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0.\end{array}\right.$ (A.2)
By straightforward calculations, this condition is the same as $\rho<\min(\rho_{1},\rho_{2})$, where
$\rho_{1}={}$ $\left[\left(\frac{-\mu_{u}}{\sigma_{u}}\right)M\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\left(1+\frac{\mu_{v}}{\mu_{u}}\right)-1\right]\frac{\sigma_{u}}{\sigma_{v}},$
$\rho_{2}={}$ $\left[\left(\frac{-\mu_{v}}{\sigma_{v}}\right)M\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\left(1+\frac{\mu_{u}}{\mu_{v}}\right)-1\right]\frac{\sigma_{v}}{\sigma_{u}}.$
Proof of Proposition 4.
Note that
$\frac{\partial g(z_{u},z_{v})}{\partial z_{u}}={}$ $p_{u}\left(-\frac{1}{\sigma_{u}}\right)\phi\left(\frac{z-\mu_{u}}{\sigma_{u}}\right)\left[(\mu_{u}+\mu_{v})+(\sigma_{u}+\rho\sigma_{v})\left(\frac{z-\mu_{u}}{\sigma_{u}}\right)\right],$
$\frac{\partial g(z_{u},z_{v})}{\partial z_{v}}={}$ $p_{v}\left(-\frac{1}{\sigma_{v}}\right)\phi\left(\frac{z-\mu_{v}}{\sigma_{v}}\right)\left[(\mu_{u}+\mu_{v})+(\rho\sigma_{u}+\sigma_{v})\left(\frac{z-\mu_{v}}{\sigma_{v}}\right)\right].$
Hence, $\nabla g(0,0)>0$ if and only if
$(\mu_{u}+\mu_{v})+(\sigma_{u}+\rho\sigma_{v})\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0\quad\mbox{and}\quad(\mu_{u}+\mu_{v})+(\rho\sigma_{u}+\sigma_{v})\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0,$
which is equivalent to
$\rho<\frac{\mu_{v}/\sigma_{v}}{\mu_{u}/\sigma_{u}}\quad\mbox{and}\quad\rho<\frac{\mu_{u}/\sigma_{u}}{\mu_{v}/\sigma_{v}}.$
If $\rho>-\min(\sigma_{v}/\sigma_{u},\sigma_{u}/\sigma_{v})$, then $\nabla g(z_{u},z_{v})=0$ has a unique solution
$z_{u}^{*}=\frac{\rho\mu_{u}/\sigma_{u}-\mu_{v}/\sigma_{v}}{\rho/\sigma_{u}+1/\sigma_{v}}\quad\mbox{and}\quad z_{v}^{*}=\frac{\rho\mu_{v}/\sigma_{v}-\mu_{u}/\sigma_{u}}{\rho/\sigma_{v}+1/\sigma_{u}};$
moreover,
$\frac{\partial^{2}g(z_{u}^{*},z_{v}^{*})}{\partial z_{u}^{2}}={}$ $p_{u}\left(-\frac{1}{\sigma_{u}}\right)\phi\left(\frac{-(\mu_{u}+\mu_{v})}{\sigma_{u}+\rho\sigma_{v}}\right)\left(1+\frac{\sigma_{v}}{\sigma_{u}}\rho\right)<0,$
$\frac{\partial^{2}g(z_{u}^{*},z_{v}^{*})}{\partial z_{v}^{2}}={}$ $p_{v}\left(-\frac{1}{\sigma_{v}}\right)\phi\left(\frac{-(\mu_{u}+\mu_{v})}{\rho\sigma_{u}+\sigma_{v}}\right)\left(1+\frac{\sigma_{u}}{\sigma_{v}}\rho\right)<0,$
$\frac{\partial^{2}g(z_{u}^{*},z_{v}^{*})}{\partial z_{u}\partial z_{v}}={}$ $\frac{\partial^{2}g(z_{u}^{*},z_{v}^{*})}{\partial z_{v}\partial z_{u}}=0.$
Thus, $\nabla^{2}g(z_{u}^{*},z_{v}^{*})$ is negative definite, implying $(z_{u}^{*},z_{v}^{*})$ is the maximizer of $g(z_{u},z_{v})$. Moreover,
$g(z_{u}^{*},z_{v}^{*})={}$ $p_{u}(\sigma_{u}+\rho\sigma_{v})\phi\left(\frac{z_{u}^{*}-\mu_{u}}{\sigma_{u}}\right)\left[\left(\frac{\mu_{u}+\mu_{v}}{\sigma_{u}+\rho\sigma_{v}}\right)M\left(\frac{z_{u}^{*}-\mu_{u}}{\sigma_{u}}\right)+1\right]$
$+p_{v}(\sigma_{u}+\rho\sigma_{v})\phi\left(\frac{z_{v}^{*}-\mu_{v}}{\sigma_{v}}\right)\left[\left(\frac{\mu_{u}+\mu_{v}}{\rho\sigma_{u}+\sigma_{v}}\right)M\left(\frac{z_{v}^{*}-\mu_{v}}{\sigma_{v}}\right)+1\right]$
[…]
$+p_{v}(\sigma_{u}+\rho\sigma_{v})\phi\left(\frac{z_{v}^{*}-\mu_{v}}{\sigma_{v}}\right)\left[-\left(\frac{z_{v}^{*}-\mu_{v}}{\sigma_{v}}\right)M\left(\frac{z_{v}^{*}-\mu_{v}}{\sigma_{v}}\right)+1\right]>0,$
where the positivity follows from Lemma 1, which implies $\alpha M(\alpha)<1$, and the assumption that $\rho>-\min(\sigma_{v}/\sigma_{u},\sigma_{u}/\sigma_{v})$, which implies $\sigma_{u}+\rho\sigma_{v}>0$ and $\rho\sigma_{m}u+\sigma_{v}>0$. ∎
Appendix B Proofs for Multivariate Normal Distribution
Lemma 3.
Let $h(z)$ be as defined in (4). Then,
$h(z)=\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\right]=n\mu\left(1-\Phi\left(\frac{z-\mu}{\sigma}\right)\right)+\sigma[1+(n-1)\rho]\phi\left(\frac{z-\mu}{\sigma}\right).$
Proof of Lemma 3.
The proof follows similar calculations to those in Lemma 2. ∎
Proof of Proposition 5.
The proof follows from Lemma 3 through direct calculations. ∎
Proof of Proposition 6.
Note that
$h^{\prime}(z)={}$ $-\frac{1}{\sigma}\phi\left(\frac{z-\mu}{\sigma}\right)\left[n\mu+[1+(n-1)\rho](z-\mu)\right]$
$h^{\prime\prime}(z)={}$ $-\frac{1}{\sigma}\phi\left(\frac{z-\mu}{\sigma}\right)\left[\left(-\frac{1}{\sigma}\right)\left(\frac{z-\mu}{\sigma}\right)\left(n\mu+[1+(n-1)\rho](z-\mu)\right)+1+(n-1)\rho\right].$
Hence, if $-1/(n-1)<\rho<1$, then $h^{\prime}(z)=0$ has a unique solution
$z^{*}=\frac{(n-1)(\rho-1)}{(n-1)\rho+1}\mu;$
moreover,
$h^{\prime\prime}(z^{*})=-\frac{1}{\sigma}\phi\left(\frac{-n\mu}{\sigma[1+(n-1)\rho]}\right)[1+(n-1)\rho]<0,$
which implies $z^{*}$ is the maximizer of $h(z)$. In addition,
$h(z^{*})={}$ $\sigma[1+(n-1)\rho)]\phi\left(\frac{z^{*}-\mu}{\sigma}\right)\left[\frac{n\mu}{\sigma[1+(n-1)\rho]}M\left(\frac{z^{*}-\mu}{\sigma}\right)+1\right]$
[…]
where the positivity follows from Lemma 1 and the assumption that $\rho>-1/(n-1)$. ∎
Appendix C Proofs for Bivariate $t$ Distribution
Lemma 4.
Suppose $(U,V)$ follows a symmetric bivariate $t$ distribution. Then,
$f(z)={}$ $2\mu\left(1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)\right)+\sigma(1+\rho)\left(\frac{\delta}{\delta-2}\right)t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right).$
Proof of Lemma 4.
Each marginal ($U$ or $V$) follows a univariate $t$ distribution with mean $\mu$, variance $\sigma^{2}$, and degrees of freedom $\delta>2$. Then, $\mathbb{E}[U|U>z]$ is the expectation of a truncated $t$ distribution with truncation interval $(z,\infty)$. Hence,
$\mathbb{E}[U|U>z]=\mu+\sigma\frac{\delta\Gamma\left(\frac{\delta+1}{2}\right)}{(\delta-1)\sqrt{\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta-1}{2}}\frac{1}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)},$ (C.1)
where $\Gamma$ denotes the gamma function; see, e.g., Ho et al., (2012).
Note that the $t$ density function $t_{\delta}(z;\mu,\sigma^{2})$ is
$t_{\delta}(z;\mu,\sigma^{2})=\frac{\Gamma\left(\frac{\delta+1}{2}\right)}{\sqrt{\sigma^{2}\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta+1}{2}}.$
Therefore,
$t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)={}$ $\frac{\Gamma(\frac{\delta-1}{2})}{\sqrt{\frac{\delta}{\delta-2}(\delta-2)\pi}\Gamma(\frac{\delta-2}{2})}\left(1+\frac{(\frac{z-\mu}{\sigma})^{2}}{\frac{\delta}{\delta-2}(\delta-2)}\right)^{-\frac{\delta-1}{2}}$
[…]
where the last step follows from the gamma function’s property $\Gamma(z+1)=z\Gamma(z)$. Hence, we can rewrite (C.1) as
$\mathbb{E}[U|U>z]=\mu+\sigma\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)}\right).$ (C.2)
The conditional distribution of $V$ given $U$ is a univariate $t$ distribution (see, e.g., Ding, (2016)). In particular, $\mathbb{E}[V|U]=\mu+\rho(U-\mu)$. Thus,
$\mathbb{E}[V|U>z]={}$ $\mathbb{E}\bigl{[}\mathbb{E}[V|U]\big{|}U>z\bigr{]}=\mu+\rho\bigl{[}\mathbb{E}[U|U>z]-\mu\bigr{]}$
[…]
Therefore,
$f(z)={}$ $\mathbb{E}[D(U+V)]$
[…]
where the third equality holds because $\mathbb{E}[U+V|U>z]=\mathbb{E}[U+V|V>z]$ (symmetry), and the last equality follows from (C.2) and (C.3). ∎
Proof of Proposition 7.
The proof follows from Lemma 2 through direct calculations. ∎
Proof of Proposition 8.
Direct calculations based on the expression of $f(z)$ in Lemma 2 yield:
$f^{\prime}(z)={}$ $-\frac{1}{\sigma}t_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)\left[2\mu+(1+\rho)(z-\mu)\right],$
where the following intermediate result is used
[…]
Hence, if $-1<\rho<1$, then $f^{\prime}(0)>0$ and $f^{\prime}(z)=0$ has a unique solution
$z^{*}=\frac{\rho-1}{\rho+1}\mu.$
Furthermore, we have
[…]
Thus,
$f^{\prime\prime}(z^{*})=-\frac{1}{\sigma}t_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)(1+\rho)<0,$
proving $z^{*}$ is the maximum of $f(z)$.
Moreover,
$f(z^{*})={}$ $2\mu\left[1-T_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)\right]\left[1+\frac{\sigma(1+\rho)}{2\mu}\left(\frac{\delta}{\delta-2}\right)\frac{t_{\delta-2}\left(\frac{z^{*}-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z^{*}-\mu}{\sigma};0,1\right)}\right]$
[…]
Let $X$ be a standard univariate $t$ distribution. Then, (C.2) implies that
$\mathbb{E}[X|X>\alpha]=\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\alpha;0,1\right)}\right),$
for all $\alpha$. Clearly, $\mathbb{E}[X|X>\alpha]>\alpha$. Hence,
$\frac{1}{\alpha}\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\alpha;0,1\right)}\right)>1.$
for all $\alpha>0$. Setting $(z^{*}-\mu)/\sigma=\alpha$, which is positive by the definition of $z^{*}$, in (C.4) and using the assumption $\mu<0$, we conclude that $f(z^{*})>0$. ∎